OpenStudy (nuccioreggie):

Marneshia walked three over eight of a mile in three over five of an hour. What equation can be used to calculate her unit rate in miles per hour? three eighths over three fifths = five eighths mile per hour three eighths over three fifths = one and three fifths miles per hour three fifths over three eights = one and three fifths miles per hour three fifths over three eights = five eighths mile per hour

OpenStudy (nuccioreggie):

@cwrw238

OpenStudy (anonymous):

what grade is this!!?! ^.^

OpenStudy (anonymous):

I hope I understand your question correctly. Lets call '\[x\]' to the unit rate miles per hour, so \[\frac{ 3 }{ 5 } . x = \frac{ 3 }{ 8 }\] \[x = \frac{ 5 }{ 8 }\]

OpenStudy (nuccioreggie):

So b